Is every subsequence of a divergent sequence is divergent ? Justify.
1
Expert's answer
2020-06-18T19:10:04-0400

Let us consider two sequences:

"\left(a_k\right)_{k=1}^{\infty}, a_k = \dfrac{1}{k}; \\
\left(b_k\right)_{k=1}^{\infty}, b_k = (-1)^k k."

We construct the sequence "(c_k)_{k=1}^{\infty}" in form of "a_1, b_1, a_2, b_2, a_3, b_3," and so on.

This sequence is divergent, because we can consider the subsequences with different limits: "c_2 = -1, c_6 = -3, c_{10} = -5, \ldots , c_{4k+2} = -(2k+1), \ldots" , which has the limit of "-\infty"

and the subsequence "c_4 = 2, c_8=4, \ldots, c_{4k}=2k, \ldots," which has the limit of "+\infty."

But subsequence "c_1, c_3, \ldots, c_{2k-1} = \dfrac{1}{k},\ldots," has the limit of 0, so it is not divergent.